# Offline Robotic World Model: Learning Robotic Policies without a Physics Simulator

### v1

@@ Abstract
Reinforcement Learning (RL) has demonstrated impressive capabilities in robotic control but remains challenging due to high sample complexity, safety concerns, and the sim-to-real gap. While offline RL eliminates the need for risky real-world exploration by learning from pre-collected data, it suffers from distribution shift, limiting policy generalization. Model-Based RL (MBRL) addresses this by leveraging predictive models for synthetic rollouts, yet existing approaches often lack robust uncertainty estimation, leading to compounding errors in offline settings. We introduce Offline Robotic World Model (RWM-O), a model-based approach that explicitly estimates epistemic uncertainty to improve policy learning without reliance on a physics simulator. By integrating these uncertainty estimates into policy optimization, our approach penalizes unreliable transitions, reducing overfitting to model errors and enhancing stability. Experimental results show that RWM-O improves generalization and safety, enabling policy learning purely from offline data and advancing scalable, data-efficient RL for robotics.
Keywords: Offline Reinforcement Learning, Model-Based Reinforcement Learning, Uncertainty Estimation
Figure 1: Overview of RWM-O and MOPO-PPO. RWM-O is trained on offline data to predict long-horizon dynamics with an ensemble-based uncertainty estimate. The shade visualizes epistemic uncertainty, which is penalized during policy optimization to avoid overfitting to unreliable predictions. Actions are omitted for clarity. Our framework enables direct policy learning without simulators, bridging the gap between model-based offline RL and real-world robotic deployment.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated significant success in robotic control, enabling policies that outperform traditional control techniques in complex tasks [1, 2, 3, 4]. However, applying RL directly to real-world robotics remains challenging due to the high sample complexity and the risks associated with exploratory interactions [5, 6]. A common alternative is training policies in high-fidelity physics simulators, which offer a safe and efficient environment for data collection and optimization [7, 8, 9]. However, simulators struggle to capture real-world dynamics with perfect accuracy, leading to the well-known sim-to-real gap, which degrades policy performance when transferred to physical hardware [10, 11]. To address these challenges, offline RL has emerged as a promising paradigm by enabling policy learning entirely from pre-collected real-world data, eliminating the need for risky online exploration [12, 13, 14, 15]. Despite its promise, offline RL is prone to distribution shift, where the learned policy encounters states that are insufficiently represented in the dataset, leading to extrapolation errors and poor generalization [13, 14, 12]. Model-free offline RL methods mitigate this by enforcing strict constraints to keep the policy within the dataset’s support, but at the cost of limiting generalization beyond observed transitions [16, 17, 18].
Model-Based Reinforcement Learning (MBRL) provides an attractive alternative by learning a predictive model of environment dynamics, allowing the generation of synthetic experience for policy optimization [19, 20]. This is particularly advantageous in offline RL for three reasons [15]. First, model-based methods benefit from denser supervision, as the dynamics model is trained on all dataset transitions, even when rewards are sparse [21]. Second, training the model via supervised learning provides more stable gradients compared to bootstrapped value estimation. Finally, well-established uncertainty estimation techniques from supervised learning can be leveraged to mitigate distribution shift [22, 23, 24, 25]. Recent work on the Robotic World Model (RWM) has demonstrated the potential of neural network-based dynamics models for policy learning [26]. RWM effectively captures long-horizon predictive dynamics without domain-specific inductive biases, enabling robust policy optimization through Model-Based Policy Optimization with Proximal Policy Optimization (MBPO-PPO) [27]. However, existing formulations lack explicit uncertainty estimation, making them susceptible to distribution shift in offline settings. Without uncertainty awareness, RWM requires online corrections to mitigate errors introduced by out-of-distribution states [12, 28, 20], which limits its applicability in fully offline RL [29, 30].
In this work, we extend RWM by incorporating uncertainty-aware dynamics modeling to enable robust policy learning in offline RL. We propose an ensemble-based architecture that explicitly estimates model uncertainty, allowing the policy to incorporate uncertainty-aware rollouts. To leverage this information, we adapt Model-Based Offline Policy Optimization (MOPO) [15] to PPO, enabling safer policy learning by penalizing high-uncertainty state transitions without requiring additional real-world interactions. Our approach enables policy learning purely from offline robotic data, eliminating the need for a physics simulator or online interactions. To our knowledge, our locomotion experiments mark the first demonstration of a locomotion policy learned entirely from offline data in a model-based fashion, with successful deployment on hardware. By enhancing the reliability and adaptability of MBRL, our approach contributes toward scalable, data-efficient policy learning, facilitating broader adoption of offline RL in real-world robotics. Supplementary videos for this work are available at https://sites.google.com/view/corl2025-rwm-o/home.
@@ 2.1 Model-Based Reinforcement Learning
Model-Based Reinforcement Learning (MBRL) constructs an explicit predictive model of environment dynamics, improving sample efficiency and facilitating long-horizon planning. Unlike model-free RL, which directly learns policies from data, MBRL generates synthetic rollouts to optimize policies, reducing reliance on real-world interactions. Early approaches such as PILCO leverage Gaussian processes to model system dynamics, achieving data-efficient learning but facing scalability issues in high-dimensional spaces [12]. Local linear models, as seen in guided policy search [28], improve tractability in complex control tasks [31]. Neural network-based dynamics models have since become the dominant approach, allowing for greater expressivity and scalability [21, 32]. One notable example is Model-Based Policy Optimization (MBPO), which interleaves model-generated rollouts with model-free updates, selectively trusting the learned model to improve stability [27].
Recent advancements in latent-space world models have further improved MBRL. PlaNet introduces a compact latent dynamics model for planning directly in an abstract space, making it computationally feasible for long-horizon tasks [19]. Dreamer extends this idea by integrating an actor-critic framework, achieving strong performance in continuous control [20, 33, 34]. A major challenge in MBRL is model fidelity and its impact on policy learning. Robotic World Model (RWM) addresses this by introducing a self-supervised, autoregressive neural simulator that learns long-horizon environment dynamics without domain-specific inductive biases [26]. Unlike previous methods, RWM enables stable policy optimization via MBPO-PPO, demonstrating successful deployment in real-world robotic systems.
@@ 2.2 Uncertainty Estimation and Distribution Shift Mitigation
Incorporating uncertainty estimation into MBRL is crucial, especially in offline settings where the lack of online interactions limits the ability to correct model inaccuracies. One of the most effective and widely used techniques for model uncertainty estimation is the use of bootstrap ensembles, which have been successfully applied in various MBRL frameworks to mitigate overestimation caused by distribution shifts in offline RL [23, 13, 15]. Probabilistic Ensembles with Trajectory Sampling (PETS) pioneers the use of probabilistic ensembles in MBRL, improving robustness by leveraging multiple network predictions for decision-making [13]. To further improve uncertainty estimation and its integration into policy learning, several advanced techniques have been developed. MOUP introduces ensemble dropout networks for uncertainty estimation while incorporating a maximum mean discrepancy constraint into policy optimization, ensuring bounded state mismatch and improving policy learning in offline settings [35]. SUMO characterizes uncertainty by measuring the cross-entropy between model dynamics and true dynamics using a k-nearest neighbor search method [36]. This technique provides an alternative perspective on estimating uncertainty beyond conventional model ensembles.
Uncertainty-aware policy optimization is key to mitigating the risks of distribution shift. MORPO explicitly quantifies the uncertainty of predicted dynamics and incorporates a reward penalty to balance optimism and conservatism in policy learning [37]. Similarly, COMBO tackles uncertainty estimation challenges by enforcing conservatism in value estimation without relying on explicit uncertainty quantification [38]. Our work extends these ideas by integrating uncertainty estimation into RWM and adapting MOPO to PPO, enhancing stability and robustness in offline robotic control.
@@ 3.1 Offline Model-Based Reinforcement Learning
We formulate the problem by modeling the environment as a Partially Observable Markov Decision Process (POMDP) [39], defined by the tuple $\left({\mathcal{S}},{\mathcal{A}},{\mathcal{O}},T,R,O,\gamma\right)$, where ${\mathcal{S}}$, ${\mathcal{A}}$, and ${\mathcal{O}}$ denote the state, action, and observation spaces, respectively. The transition kernel $T:{\mathcal{S}}\times{\mathcal{A}}\to{\mathcal{S}}$ captures the environment dynamics, while the reward function $R:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to{\mathbb{R}}$ maps transitions to scalar rewards. The agent seeks to learn a policy $\pi_{\theta}:{\mathcal{O}}\to{\mathcal{A}}$ that maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t\geq 0}\gamma^{t}r_{t}\right]$, where $r_{t}$ is the reward at time $t$ and $\gamma\in\left[0,1\right]$ is the discount factor.
In offline RL, the agent is trained using a fixed dataset of transitions collected by one or a mixture of behavior policies. Unlike online RL, the agent cannot interact with the environment, making it susceptible to extrapolation errors when encountering out-of-distribution states [29, 30]. Many offline RL methods impose conservatism by constraining the policy to remain within the dataset’s support, limiting generalization beyond observed trajectories [17, 18]. MBRL provides an alternative approach by learning a world model of the environment, which can be leveraged for policy optimization in imagination [40]. Instead of relying solely on offline data for policy learning, MBRL methods construct a predictive model of the transition dynamics, allowing the agent to generate synthetic rollouts and expand the effective training data. This enables more efficient learning and better generalization, which is particularly important in offline settings [15]. However, model errors can accumulate over long-horizon predictions, introducing bias and degrading policy performance. Addressing these challenges requires accurate long-horizon modeling and robust handling of model errors.
@@ 3.2 Robotic World Model
Robotic World Model (RWM) is a neural network-based world model designed to accurately capture long-horizon robotic dynamics without requiring domain-specific inductive biases [26]. Unlike conventional methods that rely on structured priors or handcrafted features, RWM generalizes across diverse robotic systems using a self-supervised, dual-autoregressive training framework. RWM learns to predict future observations autoregressively, feeding its own predictions back into the model to simulate long-horizon rollouts. The training objective minimizes a multi-step prediction loss
${\mathcal{L}}=\frac{1}{N}\sum_{k=1}^{N}\alpha^{k}\left[L_{o}\left(o^{\prime}_{t+k},o_{t+k}\right)+L_{c}\left(c^{\prime}_{t+k},c_{t+k}\right)\right],$ (1)
where $N$ denotes the autoregressive prediction steps, $o^{\prime}_{t+k}$ and $c^{\prime}_{t+k}$ are the predicted observations and privileged information (e.g., contact information), $L_{o}$ and $L_{c}$ measure the prediction discrepancy, and $\alpha_{k}$ is a decay factor. The predicted observation $k$ steps ahead of time $t$ can be written as
@@
where $M$ denotes the history horizon steps. Through this design, RWM achieves robust trajectory forecasting across manipulation and locomotion tasks. Policies trained with RWM have been zero-shot deployed on hardware, achieving high-fidelity control without requiring additional fine-tuning.
Model-Based Policy Optimization (MBPO) trains policies from learned world models by interleaving real environment data with simulated imagination rollouts [27]. RWM extends MBPO to PPO with extensive on-policy updates, resulting in MBPO-PPO, a model-based RL framework tailored for long-horizon robotic control [26]. MBPO-PPO follows the standard MBPO approach but benefits from high-fidelity RWM rollouts, which significantly improve stability and accuracy during policy learning.
@@ 4 Approach
4.1 Uncertainty Quantification with Offline Robotic World Model
@@
To this end, we introduce Offline Robotic World Model (RWM-O), where we explicitly incorporate uncertainty quantification into the dynamics model based on RWM. We aim to design an uncertainty estimator that captures both epistemic uncertainty, which arises due to limited training data, and aleatoric uncertainty, which reflects inherent stochasticity in the environment. Bootstrap ensembles have been shown to provide consistent estimates of uncertainty [41] and have been successfully applied in MBRL to improve robustness [13]. RWM-O extends RWM by introducing an ensemble-based uncertainty-aware architecture. Specifically, we apply bootstrap ensembles to the observation prediction head of the model after a shared recurrent feature extractor (e.g., a GRU). Each ensemble element $b$ independently predicts the mean and variance of a Gaussian distribution over the next observation
@@
During inference, the ensemble mean is used as the predicted next observation, while the variance across ensemble members provides an estimate of epistemic uncertainty $u_{p_{\phi}}$, depending on the world model $p_{\phi}$. The training overview is visualized in Fig. 1.
@@
By incorporating uncertainty-aware modeling into the dynamics learning process, RWM-O enables robust trajectory forecasting in offline settings. The explicit quantification of epistemic uncertainty allows for uncertainty-informed policy optimization, preventing the policy from over-relying on uncertain model predictions.

### v3

# Uncertainty-Aware Robotic World Model Makes Offline Model-Based Reinforcement Learning Work on Real Robots

@@ Abstract
Reinforcement Learning (RL) has achieved impressive results in robotics, yet high-performing pipelines remain highly task-specific, with little reuse of prior data. Offline Model-based RL (MBRL) offers greater data efficiency by training policies entirely from existing datasets, but suffers from compounding errors and distribution shift in long-horizon rollouts. Although existing methods have shown success in controlled simulation benchmarks, robustly applying them to the noisy, biased, and partially observed datasets typical of real-world robotics remains challenging. We present a principled pipeline for making offline MBRL effective on physical robots. Our RWM-U extends autoregressive world models with epistemic uncertainty estimation, enabling temporally consistent multi-step rollouts with uncertainty effectively propagated over long horizons. We combine RWM-U with MOPO-PPO, which adapts uncertainty-penalized policy optimization to the stable, on-policy PPO framework for real-world control. We evaluate our approach on diverse manipulation and locomotion tasks in simulation and on real quadruped and humanoid, training policies entirely from offline datasets. The resulting policies consistently outperform model-free and uncertainty-unaware model-based baselines, and fusing real-world data in model learning further yields robust policies that surpass online model-free baselines trained solely in simulation.
Figure 1: Overview of RWM-U and MOPO-PPO. RWM-U augments autoregressive world models with ensemble-based uncertainty estimation, quantifying epistemic uncertainty (shade) that arises from limited or biased offline data (left). By propagating this uncertainty consistently over long imagined rollouts, RWM-U identifies regions where predictions become unreliable under distribution shift. This uncertainty signal is then used during policy learning to penalize high-risk imagined transitions by MOPO-PPO (middle), enabling stable long-horizon rollouts and making fully offline model-based reinforcement learning practical on real robots (right).
@@ 1 Introduction
Offline reinforcement learning (RL) aims to learn policies from fixed datasets and has demonstrated the ability to extract behaviors that outperform those present in the data, making it appealing for robotic control where collected experience is often limited, expensive, or suboptimal (Deisenroth and Rasmussen, 2011; Chua et al., 2018; Clavera et al., 2018; Yu et al., 2020). Unlike online RL, which continuously interacts with the environment to collect new trajectories and correct modeling or policy errors, offline RL prohibits any additional environment interaction once the dataset is collected (Levine et al., 2020; Li et al., 2025). As a result, all components of the learning system, including dynamics modeling, long-horizon rollouts, and policy optimization, must operate entirely on a static dataset, without the corrective feedback loop afforded by online exploration (Prudencio et al., 2023). This fundamental constraint reshapes how errors accumulate and must be managed during learning, motivating algorithmic designs that remain robust under severe distributional mismatch.
In practice, such static datasets are often biased toward the behavior policy that generated them and cover a narrow subset of the state-action space (Chua et al., 2018; Clavera et al., 2018; Deisenroth and Rasmussen, 2011). These issues lead to distribution shift, where deployed policies encounter states outside the dataset’s support, making predictions unreliable. Some model-free offline RL methods mitigate this by enforcing strict constraints to keep the policy within the dataset’s support, but at the cost of limiting generalization beyond observed transitions (Agarwal et al., 2019; Kumar et al., 2019; Wu et al., 2019). In contrast, model-based RL (MBRL) tackles the problem by learning a predictive dynamics model and applying techniques such as uncertainty-aware modeling and uncertainty-penalized policy optimization to reduce the risk of unreliable predictions (Gal et al., 2016; Lakshminarayanan et al., 2017; Kuleshov et al., 2018; Ovadia et al., 2019; Yu et al., 2020). While these methods have shown success in controlled simulation benchmarks, their reliable deployment in real robotics remains challenging, as noisy, biased, and partially observed datasets demand both accurate long-horizon modeling and stable policy learning, thereby constraining the applicability of existing methods to the real world (see Sec. A.5).
In this work, we present a principled pipeline for making model-based offline RL effective in real-world robotic settings. Our Uncertainty-Aware Robotic World Model (RWM-U) extends dynamics modeling with autoregressive long-horizon prediction and ensemble-based epistemic uncertainty estimation, enabling temporally consistent multi-step rollouts. We integrate RWM-U with MOPO-PPO, an adaptation of MOPO’s uncertainty-penalized optimization to PPO, which has been widely adopted in modern real-world robotic control (Yu et al., 2020). Unlike MOPO’s one-step setup, our method must propagate and manage uncertainty over 100-step episodic rollouts, controlling compounding model errors without any further data collection. We validate our pipeline on diverse simulated tasks and real quadruped and humanoid locomotion, where policies are trained entirely offline from existing datasets and successfully deployed on ANYmal D and Unitree G1 robots. We show that fusing real-world data in model learning yields robust policies that surpass state-of-the-art online model-free baselines trained solely in simulation. To our best knowledge, this is the first demonstration of uncertainty-penalized offline MBRL operating full-scale tasks on a physical robot. Our results show that with the right integration of modeling, uncertainty handling, and policy optimization, offline RL can be a practical framework for reusing past robotic experience—effectively reusing both simulation logs and real-world data across different tasks rather than treating each development as entirely isolated. Supplementary videos for this work are available at https://sites.google.com/view/uncertainty-aware-rwm.
@@ 2.1 Model-Based Reinforcement Learning
Model-Based Reinforcement Learning (MBRL) constructs an explicit predictive model of environment dynamics, improving sample efficiency and facilitating long-horizon planning. Unlike model-free RL, which directly learns policies from data, MBRL generates synthetic rollouts to optimize policies, reducing reliance on real-world interactions. Early approaches such as PILCO leverage Gaussian processes to model system dynamics, achieving data-efficient learning but facing scalability issues in high-dimensional spaces (Deisenroth and Rasmussen, 2011). Local linear models, as seen in guided policy search (Levine and Koltun, 2013), improve tractability in complex control tasks (Kumar et al., 2016). Neural network-based dynamics models have since become the dominant approach, allowing for greater expressivity and scalability (Ebert et al., 2018; Kaiser et al., 2019). One notable example is Model-Based Policy Optimization (MBPO), which interleaves model-generated rollouts with model-free updates, selectively trusting the learned model to improve stability (Janner et al., 2019). Recent advancements in latent-space world models have further improved MBRL. PlaNet introduces a compact latent dynamics model for planning directly in an abstract space, making it computationally feasible for long-horizon tasks (Hafner et al., 2019b). Dreamer extends this idea by integrating an actor-critic framework, achieving strong performance in continuous control (Hafner et al., 2019a; 2020; 2023). Robotic World Model (RWM) introduces a self-supervised, autoregressive neural simulator that learns long-horizon environment dynamics without domain-specific inductive biases (Li et al., 2025). While these methods perform well in online settings, where model errors can be corrected through additional environment interactions, their applicability to offline settings remains underexplored.
@@ 2.2 Uncertainty Estimation and Distribution Shift Mitigation
Incorporating uncertainty estimation into MBRL is crucial, especially in offline settings where the lack of online interactions limits the ability to correct model inaccuracies. One of the most effective and widely used techniques for model uncertainty estimation is the use of bootstrap ensembles, which have been successfully applied in various MBRL frameworks to mitigate overestimation caused by distribution shifts in offline RL (Lakshminarayanan et al., 2017; Chua et al., 2018; Yu et al., 2020). Probabilistic Ensembles with Trajectory Sampling (PETS) pioneers the use of probabilistic ensembles in MBRL, improving robustness by leveraging multiple network predictions for decision-making (Chua et al., 2018). To further improve uncertainty estimation and its integration into policy learning, several advanced techniques have been developed. MOUP introduces ensemble dropout networks for uncertainty estimation while incorporating a maximum mean discrepancy constraint into policy optimization, ensuring bounded state mismatch and improving policy learning in offline settings (Zhu et al., 2024). SUMO characterizes uncertainty by measuring the cross-entropy between model dynamics and true dynamics using a k-nearest neighbor search method (Qiao et al., 2024). This technique provides an alternative perspective on estimating uncertainty beyond conventional model ensembles. Uncertainty-aware policy optimization is key to mitigating the risks of distribution shift. MORPO explicitly quantifies the uncertainty of predicted dynamics and incorporates a reward penalty to balance optimism and conservatism in policy learning (Guo et al., 2022). Similarly, CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2021), and COMBO (Yu et al., 2021) tackle uncertainty estimation challenges by enforcing conservatism in value estimation without relying on explicit uncertainty quantification. While these methods achieve impressive performance in controlled simulation benchmarks, applying them to real-world robotics remains a significant hurdle, where reliability and robustness demand both accurate long-horizon modeling and stable policy learning. We highlight some of these real challenges in Sec. A.5.
@@ 3.1 Offline Model-Based Reinforcement Learning
We formulate the problem by modeling the environment as a Partially Observable Markov Decision Process (POMDP) (Sutton and Barto, 2018), defined by the tuple $\left({\mathcal{S}},{\mathcal{A}},{\mathcal{O}},T,R,O,\gamma\right)$, where $\mathcal{S}$}, $\mathcal{A}$}, and $\mathcal{O}$} denote the state, action, and observation spaces, respectively. The transition kernel $T:{\mathcal{S}}\times{\mathcal{A}}\to{\mathcal{S}}$ captures the environment dynamics, while the reward function $R:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to{\mathbb{R}}$ maps transitions to scalar rewards. The agent seeks to learn a policy $\pi_{\theta}:{\mathcal{O}}\to{\mathcal{A}}$ that maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t\geq 0}\gamma^{t}r_{t}\right]$, where $r_{t}$ is the reward at time $t$ and $\gamma\in\left[0,1\right]$ is the discount factor.
In offline RL, the agent is trained using a fixed dataset of transitions collected by one or a mixture of behavior policies. Unlike online RL, the agent cannot interact with the environment, making it susceptible to extrapolation errors when encountering out-of-distribution states (Fujimoto et al., 2019; Fu et al., 2020). Many offline RL methods impose conservatism by constraining the policy to remain within the dataset’s support, limiting generalization beyond observed trajectories (Kumar et al., 2019; Wu et al., 2019). MBRL provides an alternative approach by learning a world model of the environment, which can be leveraged for policy optimization in imagination (Sutton, 1991). Instead of relying solely on offline data for policy learning, MBRL methods construct a predictive model of the transition dynamics, allowing the agent to generate synthetic rollouts and expand the effective training data. This enables more efficient learning and better generalization, which is particularly important in offline settings (Yu et al., 2020). However, model errors can accumulate over long-horizon predictions, introducing bias and degrading policy performance. Addressing these challenges requires accurate long-horizon modeling and robust handling of model errors.
@@ 3.2 Robotic World Model
Robotic World Model (RWM) is a neural network-based world model designed to accurately capture long-horizon robotic dynamics without requiring domain-specific inductive biases (Li et al., 2025). Unlike conventional methods that rely on structured priors or handcrafted features, RWM generalizes across diverse robotic systems using a self-supervised, dual-autoregressive training framework. RWM learns to predict future observations autoregressively, feeding its own predictions back into the model to simulate long-horizon rollouts. The training objective minimizes a multi-step prediction loss
${\mathcal{L}}=\frac{1}{N}\sum_{k=1}^{N}\alpha^{k}\left[L_{o}\left(o^{\prime}_{t+k},o_{t+k}\right)\right],$ (1)
where $N$ denotes the autoregressive prediction steps, $o^{\prime}_{t+k}$ is the predicted observations, $L_{o}$ measures the prediction discrepancy, and $\alpha_{k}$ is a decay factor. The predicted observation $k$ steps ahead of time $t$ can be written as
@@
where $M$ denotes the history horizon steps. Through this design, RWM achieves robust trajectory forecasting across manipulation and locomotion tasks. Policies trained with RWM have been zero-shot deployed on hardware, achieving high-fidelity control without requiring additional fine-tuning. However, RWM lacks uncertainty estimation, requiring policy training to be conducted online, where additional interactions induced by the current policy are leveraged to correct model errors, making this approach infeasible in offline settings.
@@ 4 Approach
4.1 Uncertainty Quantification with Uncertainty-Aware Robotic World Model
@@
To this end, we introduce Uncertainty-Aware Robotic World Model (RWM-U), where we explicitly incorporate uncertainty quantification into the dynamics model based on RWM. We aim to design an uncertainty estimator that captures both epistemic uncertainty, which arises due to limited training data, and aleatoric uncertainty, which reflects inherent stochasticity in the environment. Bootstrap ensembles have been shown to provide consistent estimates of uncertainty (Bickel and Freedman, 1981) and have been successfully applied in MBRL to improve robustness (Chua et al., 2018). RWM-U extends RWM by introducing an ensemble-based uncertainty-aware architecture. Specifically, we apply bootstrap ensembles to the observation prediction head of the model after a shared recurrent feature extractor (e.g., a GRU). Each ensemble element $b$ independently predicts the mean and variance of a Gaussian distribution over the next observation
@@
During inference, the ensemble mean is used as the predicted next observation, while the variance across ensemble members provides an estimate of epistemic uncertainty $u_{p_{\phi}}$, depending on the world model $p_{\phi}$. The system overview is visualized in Fig. 1 and the training pipeline is illustrated in Fig. 2. We provide more details on model architecture and hyperparameters in Sec. A.2.
@@
By incorporating uncertainty-aware modeling into the dynamics learning process, RWM-U enables robust trajectory forecasting in offline settings with uncertainty effectively propagated over long horizons. The explicit quantification of epistemic uncertainty allows for uncertainty-informed policy optimization, preventing the policy from over-relying on uncertain model predictions.
Figure 2: Training diagram of RWM-U and MOPO-PPO. RWM-U extends autoregressive world models with ensemble-based uncertainty estimation. Each ensemble member independently predicts a Gaussian distribution over the next observation. The variance within each prediction captures aleatoric uncertainty, while the variance across ensemble means estimates epistemic uncertainty arising from limited or biased offline data. During policy training in imagination, MOPO-PPO penalizes high-risk imagined transitions using this uncertainty signal, balancing task performance against model confidence.